Calculate the ratio of the critical angle of diamond to the critical angle of crown glass if both are surrounded by a vacuum.




1
Expert's answer
2021-11-14T17:15:28-0500

"\theta_d=\sin^{-1}(1/n_d)=\sin^{-1}(1/2.417)=24.44°"


"\theta_c=\sin^{-1}(1/n_c)=\sin^{-1}(1/2.1.75)=34.85°"


"\theta_d/\theta_c=24.44°/34.85°=0.7" . Answer